# A quick probability-oriented introduction to operator splitting methods

### v1

@@ 4. The Trotter-Kato formula
Theorem 6 can be formulated for a finite number of semigroups (e.g. [280]).
@@
This result belongs to H.E. Trotter and T. Kato. Yu.L. Daletskii also obtained similar product formulae at the same time (see [93] for references).
@@
The choice of the topology matters: the convergence may not hold in a stronger topology [265].
@@
The stability assumption (4.1) cannot be dropped [221].
@@
in some Hilbert space. The operator $H$ is essentially self-adjoint (that is, has a self-adjoint extension) under some rather mild assumptions on $V,$ so the semigroup generated by $iH$ is an unitary $C_{0}-$semigroup, in particular [^12] We skip all technical moments concerning to operators not being closed, the existence of multiple extensions and do not mention symmetric and skew-adjoint operators even though treating these issues is an integral part of quantum mechanics.. Since the semigroups for $\frac{i}{2}\Delta$ and $iV$ are
@@
which is an polygonal approximation for the Feynman integral over all histories (trajectories in the space $C_{x}([0;t])$ of continuous trajectories starting at $x$[289, 290, 151, 205]
@@
([289, 308, 205, 240, 171, 290]) It is customary for textbooks on quantum mechanics to derive the Feynman-Kac formula for Shrödinger operators as a corollary of the Trotter-Kato formula. To understand the idea formally recall that the solution to the following autonomous homogeneous PDE (both assumptions are used in what follows)
@@
It should be noted that the precise formulations of physical and probabilistic versions of the Feynman-Kac formula may be dictated by different points of focus: physicists are often more concerned with results for the Laplacian and singular or irregular (in different senses) potentials (e.g. those belonging to Kato classes or $L_{\infty}({\mathbb{R}}^{d})$) [^13] On this path, various approximations of the original problem with ACPs with ”better” coefficients are needed along with the corresponding limit theorems.. This remark applies primarily to textbooks.
@@
([229]; cf. Example 13) It is possible to derive the Girsanov theorem using the Trotter-Kato formula (under some assumptions on the regularity of drift)141414[229] also uses the Euler-Maruyama scheme to prove the Girsanov theorem.. Following the original source, we explain some intuition behind this. Let $(T_{t})_{t\geq 0}$ be the heat semigroup. The semigroup $(S_{t})_{t\geq 0}$
@@
The Trotter-Kato formula for self-adjoint operators [289, 171] is usually written [^16] Here $(\mathrm{e}^{tC})_{t\geq 0}$ denote the semigroup with generator $C$. as
@@
The importance of these versions for physics follows from the Stone theorem: on Hilbert spaces, every unitary $C_{0}-$group has a generator of the form $iA$ for some self-adjoint $A$ and every symmetric $C_{0}-$semigroup has a generator of the form $-A$ for some bounded below self-adjoint $A,$ and this is exactly the situation of quantum mechanics. The Trotter-Kato formula admits special proofs independent of the Trotter-Kato theorems in this case. In particular, if $A_{1}+A_{2}$ is densely defined and self-adjoint on $D(A_{1})\cap D(A_{2})$ the proof is rather direct. For further references and extensions see [240, 205, 308], in particular. However, the rate of the convergence was found only recently and under additional assumptions (see e.g. Example 27 and related references).
@@
To formulate a formal non-autonomous version of the Trotter-Kato formula, consider operators $A_{0},A_{1},A_{2}$ such that the corresponding non-autonomous ACPs are well-posed with propagators $(T^{(k)}_{s,t})_{0\leq s\leq t\leq T},k=0,1,2,$ respectively. Additionally, assume that $D(A_{0}(t)),D(A_{1}(t)),D(A_{2}(t)),t\in[0;T],$ do not depend on $t,$ and $A_{0}(t)=A_{1}(t)+A_{2}(t)$ on $D(A_{0}(0)),t\in[0;T].$ Then one expects for any $s,t\in[0;T],s\leq t,$
@@
Note that the condition of the domains being constant is a rather limiting one and excludes time-dependent boundary conditions in general as BCs are usually incorporated into the very description of a domain.
@@
([121]; also [25]) The Trotter-Kato formula can be used to derive an approximation of a non-autonomous ACP by autonomous ACPs with constant coefficients (cf. Example 18). For that, assume that $(U_{s,t})_{0\leq s\leq t\leq T}$ is the propagator for some well-posed non-autonomous ACP and define semigroups $(T^{(s,n,k)}_{l})_{s\leq l\leq T},k=\overline{1,n},n\in{\mathbb{N}},s\in[0;T],$ for the equations
@@
In Example 14, $D(\Delta)\cap D(V\cdot)$ may be too small or just $\{0\}$ even if the sum $\frac{1}{2}\Delta+V$ can be defined via quadratic forms [205, Chapter 10][171, Chapter 9]. In fact, one can use the Trotter-Kato formula to define the generalized sum of two operators with incompatible domains [151, Remark 8.17].
@@
The original formulation of the Trotter-Kato theorem for PDEs require boundary conditions to be time-independent. In general, adding non-autonomous boundary and initial conditions leads to significant technical issues.
@@
The condition for the sets $D$ and $(\lambda\mathrm{Id}-A)$ to be dense in $X$ implies (in a nontrivial way) that a densely defined closed $A$ exists and is indeed a generator of a $C_{0}-$semigroup so this condition can be seen [121] as an extension of the Hille-Yosida theorem171717the implications $(1)\Rightarrow(2),(3)$ in Theorem 4. In fact, some formulations of the Trotter-Kato formula just require $A+B$ to be a generator of a $C_{0}-$semigroup. The same applies to the Chernoff product formula (e.g. [94]). In practical terms, this means that one has to be careful with domains when trying to replicate the original proof.
@@
This is the second part of the proof. Since $A_{n}\to A,n\to\infty,$ the Trotter-Kato theorem implies (4.5). But there are no possibilities to either have a bound for this expression or to replace strong convergence with convergence in the operator norm in general because the proof of the Trotter-Kato theorem does not produce such results internally. As a result, the standard Chernoff product formula and Trotter-Kato formula are formulated in the terms of strong convergence and with no bounds, which may not be ideal for applications.
It may be possible to refine (4.4) and/or strengthen the convergence in the Trotter-Kato theorem for special classes of generators (e.g. self-adjoint or maximal accretive ones). Then one can combine improved estimates in (4.4) (in particular, those obtained by using properties of the Poisson distribution) and consistency estimates for the Trotter-Kato theorem to yield the conclusion about the convergence in the operator norm and to provide estimates for its speed [68, 344, 279, 343, 345] (see also [280, Chapter 3, Lemma 3.5]). This may be compared with [154, 153] where estimates of the speed of convergence often include an original element $x.$ See also [200, 201] for other results of this type. Some other references concerning bounds for the rate of convergence will be added later.
@@
This principle extends to other settings and situations as soon as an approximation scheme arises (see [200] for the discussion in the setting of evolution equations and [148, Chapter 3] for the discussion in the case of the Trotter-Kato formula).
Conditions (3.5), (3.6) and (4.1) give stability of a scheme while the convergence of generators or the behavior of the family $(V(t))_{t\geq 0}$ at 00 are consistency conditions in our case etc. However, one should not forget that, contrary to the case of the Lax-Richtmyer theorem, the principle ”stability and consistency imply convergence” is no longer a rigorous statement and does not also reflect the full nature of proofs since such proofs control ranges and domains of unbounded operators and ensure the possibility for an operator to be a generator of $C_{0}-$semigroup. On other side, this principle remains a extremely powerful tool often encountered in situations where various bounds are being developed.
@@
For a discussion of consistency in the terms of the resolvent see [201]; for a discussion of a possible trade-off between consistency and stability, [201, 189]; for a version for inhomogeneous ACPs, [189]181818[300] gives an example of a non-stable additive splitting scheme..
@@
Compositions of generators of contraction semigroups automatically are stable; alternatively, a method is expected to be stable if the operators involved commute [37]. This is a common duality in the theory of (multiplicative) splitting methods.
@@
Let $(R_{s,t})_{0\leq s\leq t}$ be the solution operator of an ACP with initial value $y_{0}$ and let $(\hat{R}_{k,n})_{k=\overline{0,n-1},n\in{\mathbb{N}}}$ be the solution operator of some approximation scheme on uniform mesh:
$\hat{y}_{n,k+1}$ $=\hat{R}_{k,n}\hat{y}_{n,k},\quad k=\overline{0,n-1},n\in{\mathbb{N}}.$
Setting $\prod_{j=\overline{n,n-1}}\hat{R}_{j,n}=\mathrm{Id}$ we have
@@
where $g\in C^{2}({\mathbb{R}})$ with bounded derivatives and such that $g(0)=0.$ Let $(T_{t})_{t\geq 0}$ be the heat semigroup and
@@
When combined with energy stability bounds for the original semigroup and Example 20, this implies that the Trotter-Kato formula is first order accurate. Note that the rate of the convergence is obtained by using probabilistic techniques.
@@
[123] also establishes $L_{p}-$estimates (in the linear case) and results for the $F_{t}\circ T_{t}-$version of the Trotter-Kato formula and for the case when one step approximation is $F_{t/2}\circ T_{t}\circ F_{t/2}$ (the Strang splitting).
For other generalizations of the Trotter-Kato, bounds, further developments, extensions to other types of semigroups and references see, besides the sources listed above, [291, 68, 260, 267, 194, 195, 205, 190, 66, 124, 262, 342, 175, 176, 268, 191, 198, 197, 128, 37, 328, 264, 330, 331, 261, 265, 67, 266, 329, 87] (both for the Chernoff product formula and the Trotter-Kato formula).
In particular, the non-autonomous case is studied in [93, 195, 262, 176, 25, 121, 128, 328, 330, 331, 266, 329]. In particular, [330, 331, 329] study time-dependent domains.
Versions of the (non-autonomous) Trotter-Kato formula in alternative settings are established in [323, 93, 119, 200, 25]. In particular, [323, 119] use the framework of Gelfand triples (rigged Hilbert spaces).
For results on nonlinear Trotter-Kato formulae, see references in [289, Supplement VIII.8].
@@
Splitting on the level of semigroups means splitting in time. Numerical schemes also include spatial discretization in practice. For results about combining time splitting and spatial discretization (in particular, a version of the Chernoff formula) see [200, 27, 26]. See also Remark 41.
@@
Revisiting examples for Feller processes, one can see that the action of the corresponding semigroup is defined and well behaved for $f\in C_{0}({\mathbb{R}}^{n})$ ($L_{p}({\mathbb{R}}^{n})$) and not necessarily for functions outside of such classes though such function may appear in applications. For one extension of splitting schemes to larger classes of functions with preservation of the rate of the convergence for SPDEs see [110].
@@
The previous sections contain a rigorous description of the Trotter-Kato formula for semigroups. Henceforward we care more about ideas and illustrations and no longer aim for precise formulations as we step into the domain of general operator splitting methods as numerical and theoretical schemes to approximate an original solution without any a priori expectations about convergence or unified results.
Abstract initial value problems (IVPs) [^20] We do not call them abstract Cauchy problem as a discussion of semigroups involved is neither provided nor needed. with no assumptions on the corresponding operators (e.g. they can be multivalued or discontinuous) are used. We mostly follow the exposition in [148, 188].
@@
In this section we write $Au$ and similar terms in the LHS of an IVP. Later in the text we will resume writing them in the RHS.
Assume $T$ is fixed. Let $h=\frac{T}{N}$ be a fixed time step for some integer $N.$ Set $t_{n}=nh$ so $\{t_{0},\ldots,t_{N}\}$ form a partition of $[0;T].$
The following scheme is a general version of the Trotter-Kato formula and is called the Lie-Trotter splitting scheme. In particular, the general version of Theorem 1 is
@@
$\frac{dv_{j,n}(t)}{dt}+A_{j}v_{j,n}=0,\quad t\in[t_{n};t_{n+1}],$
@@
$v_{m,n}(t_{n+1})\approx u(t_{n+1}),\quad n=\overline{1,N}.$
@@
$v_{1,n}(t_{n})=\widetilde{v}_{1,n-1}\Big{(}t_{n-1}+\frac{h}{2}\Big{)},\ v_{2,n}(t_{n})=v_{1,n}\Big{(}t_{n}+\frac{h}{2}\Big{)},\ \widetilde{v}_{1,n}(t_{n})=v_{2,n}(t_{n+1}),$
$v_{1,0}(0)=u_{0}.$
@@
$u(t_{n+1})\approx\widetilde{v}_{1,n}\Big{(}t_{n}+\frac{h}{2}\Big{)},\quad n=\overline{1,N}.$
One does expect the second order in the general case but, as earlier, this is not guaranteed. We will not repeat this universal remark on the absence of a universal rate for any given splitting scheme in future.
@@
The Strang splitting is quite popular not only due to it being of the second order, but also because it does not require doubling the number of computations (compared to the Lie splitting), as we have already seen in the finite-dimensional case. Indeed, if the propagator for the semigroup of $A_{k}$ is $(S^{(k)}_{t})_{t},k=\overline{1,2},$ we get
$u(T)$ $\approx\widetilde{v}_{2,N-1}\Big{(}T-\frac{h}{2}\Big{)}$
$=S^{(2)}_{t_{N-1},t_{N-1}+h/2}\circ\ldots S^{(2)}_{t_{n}+h/2,t_{n+1}}\circ S^{(1)}_{t_{n},t_{n+1}}\circ S^{(2)}_{t_{n},t_{n}+h/2}\ldots\circ S^{(2)}_{0,h/2}u_{0}$
$=S^{(2)}_{t_{N-1},t_{N-1}+h/2}\circ\ldots S^{(2)}_{t_{n-1}+h/2,t_{n}+h/2}\circ S^{(1)}_{t_{n-1},t_{n}}\circ S^{(2)}_{t_{n-2}+h/2,t_{n-1}+h/2}\circ\ldots\circ S^{(2)}_{0,h/2}u_{0}.$
The Lie-Trotter and Strang splitting schemes are multiplicative splitting schemes: the formal semigroups for $A_{1},\ldots,A_{m}$ are combined through composition exclusively.
@@
or by using the implicit backward Euler methods
@@
The Douglas-Rachford splitting can be extended to the case $m>2$[148].
@@
If the operator $A_{2}$ is computationally problematic, both Peaceman-Rachford and Douglas-Rachford splittings can be modified to avoid using it for solving the second subproblem. For instance, solving the first equation for $A_{2}(\widetilde{t}_{n},\widetilde{u}_{n})$ we get for the Peaceman-Rachford splitting
@@
There are other common additive operator splitting schemes such as the Tseng splitting (e.g. [314, 11]) and the Davis-Yin splitting (e.g. [14, 75]).
@@
Let us recall some very basic facts about this last assumption and optimization theory [28, 347] and briefly explain some terminology that is often encountered in the context of operator splitting methods in optimization. Let $B$ be a Banach space. A possibly nonlinear and multivalued operator $A\colon B\mapsto B^{*}$ is called monotone if
@@
Operator monotonicity is one of the cornerstones of optimization theory in general and convex optimization in particular. The Browder-Minty theorem states that if $A$ is additionally hemicontinuous and coercive [^21] Recall that these are also standard assumptions in the theory of SPDEs, that is, the mappings
@@
Moreover, popular and effective alternating direction methods of (Lagrange) multipliers (ADDM) are actually the reformulated Peaceman-Rachford and Douglas-Rachford splittings or other additive operator splitting schemes.
@@
([28]) A so-called forward-backward splitting can be derived as follows: we want to find $x$ such that
@@
(5.4) $\frac{du_{n+1}(t)}{dt}=B_{h}u_{n+1},\quad u_{n+1}(0)=u_{n},$
where the matrix $B$ depends on the grid and the size $h$ of the grid explicitly. If both space and time discretizations are used, we have a high-dimensional nonlinear equation.

### v2

@@ 4. The Trotter-Kato formula
Theorem 6 can be formulated for a finite number of semigroups (e.g. [Pazy12Semigroups]).
@@
This result belongs to H.E. Trotter and T. Kato. Yu.L. Daletskii also obtained similar product formulae at the same time (see [DaFo91measures] for references).
@@
The choice of the topology matters: the convergence may not hold in a stronger topology [NeidSteZa18remarks].
@@
The stability assumption (4.1) cannot be dropped [KuhnWa00Lie].
@@
in some Hilbert space. The operator $H$ is essentially self-adjoint (that is, has a self-adjoint extension) under some rather mild assumptions on $V,$ so the semigroup generated by $iH$ is an unitary $C_{0}-$semigroup, in particular [^12] We skip all technical aspects associated with the existence of multiple extensions etc., and we do not mention symmetric and skew-adjoint operators even though treating these issues is an integral part of quantum mechanics.. Since the semigroups for $\frac{i}{2}\Delta$ and $iV$ are
@@
which is an polygonal approximation for the Feynman integral over all histories (trajectories in the space $C_{x}([0;t])$ of continuous trajectories starting at $x$[ReedSai72methods, ReedSi75methods_2, Goldstein85semigroups, JohnLa00Feynman]
@@
([ReedSai72methods, Si05functional, JohnLa00Feynman, LoHiBetz20FeynmanKac, Hall13quantum, ReedSi75methods_2]) It is customary for textbooks on quantum mechanics to derive the Feynman-Kac formula for Shrödinger operators as a corollary of the Trotter-Kato formula. To understand the idea formally recall that the solution to the following autonomous homogeneous PDE (both assumptions are used in what follows)
@@
It should be noted that precise formulations of physical and probabilistic versions of the Feynman-Kac formula may be dictated by different points of focus: physicists are often more concerned with results for the Laplacian and singular or irregular (in different senses) potentials (e.g. those belonging to Kato classes or $L_{\infty}({\mathbb{R}}^{d})$) [^13] On this path, various approximations of the original problem with ACPs with ”better” coefficients are needed along with the corresponding limit theorems.. This remark applies primarily to textbooks.
@@
([Lejay18Girsanov]; cf. Example 13) It is possible to derive the Girsanov theorem using the Trotter-Kato formula (under some assumptions on the regularity of drift)141414[Lejay18Girsanov] also uses the Euler-Maruyama scheme to prove the Girsanov theorem.. Following the original source, we explain some intuition behind this. Let $(T_{t})_{t\geq 0}$ be the heat semigroup. The semigroup $(S_{t})_{t\geq 0}$
@@
The Trotter-Kato formula for self-adjoint operators [ReedSai72methods, Hall13quantum] is usually written [^16] Here $(\mathrm{e}^{tC})_{t\geq 0}$ denote the semigroup with generator $C$. as
@@
The importance of these versions for physics follows from the Stone theorem: on Hilbert spaces, every unitary $C_{0}-$group has a generator of the form $iA$ for some self-adjoint $A$ and every symmetric $C_{0}-$semigroup has a generator of the form $-A$ for some bounded below self-adjoint $A,$ and this is exactly the situation of quantum mechanics. The Trotter-Kato formula admits special proofs independent of the Trotter-Kato theorems in this case. In particular, if $A_{1}+A_{2}$ is densely defined and self-adjoint on $D(A_{1})\cap D(A_{2})$ the proof is rather direct. For further references and extensions see [LoHiBetz20FeynmanKac, JohnLa00Feynman, Si05functional], in particular. However, the rate of the convergence was found only recently and under additional assumptions (see e.g. Example 27 and related references).
@@
To formulate a formal non-autonomous version of the Trotter-Kato formula, consider operators $A_{0},A_{1},A_{2}$ such that the corresponding non-autonomous ACPs are well-posed with propagators $(T^{(k)}_{s,t})_{0\leq s\leq t\leq T},k=0,1,2,$ respectively. Additionally, assume that $D(A_{0}(t)),$ $D(A_{1}(t)),$ $D(A_{2}(t)),$ do not depend on $t,$ and $A_{0}(t)=A_{1}(t)+A_{2}(t)$ on $D(A_{0}(0)),t\in[0;T].$ Then one expects for any $s,t\in[0;T],s\leq t,$
@@
Note that the condition of the domains being constant is a rather limiting one and excludes time-dependent boundary conditions in general as BC is usually incorporated into the very description of a domain.
@@
([EnNa00One-parameter]; also [BatCsoFarNi11operator]) The Trotter-Kato formula can be used to derive an approximation of a non-autonomous ACP by autonomous ACPs with constant coefficients (cf. Example 18). For that, assume that $(U_{s,t})_{0\leq s\leq t\leq T}$ is the propagator for some well-posed non-autonomous ACP and define semigroups $(T^{(s,n,k)}_{l})_{s\leq l\leq T},k=\overline{1,n},n\in{\mathbb{N}},s\in[0;T],$ for the equations
@@
In Example 14, $D(\Delta)\cap D(V\cdot)$ may be too small or just $\{0\}$ even if the sum $\frac{1}{2}\Delta+V$ can be defined via quadratic forms [JohnLa00Feynman, Chapter 10][Hall13quantum, Chapter 9]. In fact, one can use the Trotter-Kato formula to define the generalized sum of two operators with incompatible domains [Goldstein85semigroups, Remark 8.17].
@@
The original formulation of the Trotter-Kato theorem for PDEs requires boundary conditions to be time-independent. In general, adding non-autonomous boundary and initial conditions can lead to significant technical issues.
@@
The condition for the sets $D$ and $(\lambda\mathrm{Id}-A)$ to be dense in $X$ implies (in a nontrivial way) that a densely defined closed $A$ exists and is indeed a generator of a $C_{0}-$semigroup so this condition can be seen [EnNa00One-parameter] as an extension of the Hille-Yosida theorem171717the implications $(1)\Rightarrow(2),(3)$ in Theorem 4. In fact, some formulations of the Trotter-Kato formula just require $A+B$ to be a generator of a $C_{0}-$semigroup. The same applies to the Chernoff product formula (e.g. [Da0One-parameter]). In practical terms, this means that one has to be careful with domains when trying to replicate the original proof.
@@
This is the second part of the proof. Since $A_{n}\to A,n\to\infty,$ the Trotter-Kato theorem implies (4.5). But there are no possibilities to either have a bound for this expression or to replace strong convergence with convergence in the operator norm in general because the proof of the Trotter-Kato theorem does not produce such results internally. As a result, the standard Chernoff product formula and Trotter-Kato formula are formulated in the terms of strong convergence and with no bounds.
It may be possible to refine (4.4) and/or strengthen the convergence in the Trotter-Kato theorem for special classes of generators (e.g. self-adjoint or maximal accretive ones). Then one can combine improved estimates in (4.4) (in particular, those obtained by using properties of the Poisson distribution) and consistency estimates for the Trotter-Kato theorem to yield the conclusion about the convergence in the operator norm and to provide estimates for its speed [CaZag01operator, Zag22notes, Pau04operator-norm, Zag17comments, Za20notes_chernoff_formula] (see also [Pazy12Semigroups, Chapter 3, Lemma 3.5]). This may be compared with [GoTo14convergence, GoKoTo19general] where estimates of the speed of convergence often include an original element $x.$ See also [ItoKa02evolution, ItoKa98Trotter] for other results of this type. Some other references concerning bounds for the rate of convergence will be added later.
@@
This principle extends to other settings and situations as soon as an approximation scheme arises (see [ItoKa02evolution] for a discussion in the setting of evolution equations and [GlowOshYin17Splitting, Chapter 3] for a discussion in the case of the Trotter-Kato formula).
Conditions (3.5), (3.6) and (4.1) give stability of a scheme while the convergence of generators or the behavior of the family $(V(t))_{t\geq 0}$ at 00 are consistency conditions in our case etc. However, one should not forget that, contrary to the case of the Lax-Richtmyer theorem, the principle ”stability and consistency imply convergence” is no longer a rigorous statement and does not also reflect the full nature of proofs since such proofs deal with ranges and domain of unbounded operators to obtain the generator of a $C_{0}-$semigroup. On other side, this principle remains a extremely powerful tool often encountered in situations where various bounds are being developed.
@@
For a discussion of consistency in the terms of the resolvent see [ItoKa98Trotter]; for a discussion of a possible trade-off between consistency and stability, [ItoKa98Trotter, HuItoKa01aspects]; for a version for inhomogeneous ACPs, [HuItoKa01aspects]181818[Schat99stability] gives an example of a non-stable additive splitting scheme..
@@
Compositions of generators of contraction semigroups automatically are stable; alternatively, a method is expected to be stable if the operators involved commute [Bjor98operator]. This is a common duality in the theory of (multiplicative) splitting methods.
@@
Let $(R_{s,t})_{0\leq s\leq t}$ be the solution operator of an ACP with initial value $y_{0}$ and let $(\hat{R}_{k,n})_{k=\overline{0,n-1},n\in{\mathbb{N}}}$ be solution operators of some recurrent approximation scheme on the uniform (in time) mesh:
$\hat{y}_{n,k+1}$ $=\hat{R}_{k,n}\hat{y}_{n,k},\quad k=\overline{0,n-1},n\in{\mathbb{N}},$
where $\hat{y}_{n,n},n\in{\mathbb{N}},$ converge to the original solution $R_{0,t}y_{0}.$ Setting $\prod_{j=\overline{n,n-1}}\hat{R}_{j,n}=\mathrm{Id}$ we have
@@
where $g\in C^{2}({\mathbb{R}})$ has bounded derivatives and satisfies $g(0)=0.$ Let $(T_{t})_{t\geq 0}$ be the heat semigroup and
@@
When combined with energy stability estimates for the original semigroup and Example 20, this implies that the Trotter-Kato formula is first order accurate. Note that the rate of the convergence is obtained by using probabilistic techniques.
@@
[Faou09Analysis] also establishes $L_{p}-$estimates (in the linear case) and results for the $F_{t}\circ T_{t}-$version of the Trotter-Kato formula and for the case when one step approximation is $F_{t/2}\circ T_{t}\circ F_{t/2}$ (the Strang splitting).
For other generalizations of the Trotter-Kato, bounds, further developments, extensions to other types of semigroups and references see, besides the sources listed above, [Ro93error, CaZag01operator, NeidZag98error_estimate, NeiZag99fractional, IchiTa97error, IchiTa98error, JohnLa00Feynman, Ichi04recent, CaNeidZa01accretive, FaouOsSchratz15analysis, NeidZa20convergence, Za05Trotter, HanOs10dimension, HanOs09dimension, NeidZa99fractional, IchiNeidZa04Trotter, IchiTaTaZa01note, IchiTa01norm, Fa67product, Bjor98operator, Vui10generalization, NeidSteZa18operator, VuiWresZa09general, VuiWresZa08Trotter, NeidZa99Trotter, NeidSteZa18remarks, CaNeidZa02comments, NeidSteZa19Trotter, VuiWres11product, CsoEhrFas21operator] (both for the Chernoff product formula and the Trotter-Kato formula).
In particular, the non-autonomous case is studied in [DaFo91measures, IchiTa98error, NeidZa20convergence, HanOs09dimension, BatCsoFarNi11operator, EnNa00One-parameter, Fa67product, Vui10generalization, VuiWresZa09general, VuiWresZa08Trotter, NeidSteZa19Trotter, VuiWres11product]. In particular, [VuiWresZa09general, VuiWresZa08Trotter, VuiWres11product] study time-dependent domains.
Versions of the (non-autonomous) Trotter-Kato formula in alternative settings are established in [Te68stabilite, DaFo91measures, EisenHan22variational, ItoKa02evolution, BatCsoFarNi11operator]. In particular, [Te68stabilite, EisenHan22variational] use the framework of Gelfand triples (rigged Hilbert spaces).
For results on nonlinear Trotter-Kato formulae, see references in [ReedSai72methods, Supplement VIII.8].
@@
Splitting on the level of semigroups means splitting in time. Numerical schemes also include spatial discretization in practice. For results about combining time splitting and spatial discretization (in particular, a version of the Chernoff formula) see [ItoKa02evolution, BatCsoNi09operator_splittings, BaCsoFarNi12operator_spatial]. See also Remark 42.
@@
Revisiting examples for Feller processes, one can see that the action of the corresponding semigroup is defined and well behaved for $f\in C_{0}({\mathbb{R}}^{n})$ ($L_{p}({\mathbb{R}}^{n})$) and not necessarily for functions outside of such classes though such function may appear in applications. See [DoerTeich10semigroup] for one extension of splitting schemes to larger classes of functions with preservation of the rate of the convergence for SPDEs.
@@
The previous sections contain a rigorous description of the Trotter-Kato formula for semigroups. Henceforward we care more about ideas and illustrations and no longer aim to give precise formulations as we step into the domain of general operator splitting methods as numerical and theoretical schemes to approximate an original solution without any a priori expectations about convergence or a unified exposition.
Abstract initial value problems (IVPs) [^20] We do not call them abstract Cauchy problem as a discussion of semigroups involved is neither provided nor needed. with no assumptions on the corresponding operators (e.g. they can be multivalued or discontinuous) are used. We mostly follow the exposition in [GlowOshYin17Splitting, HundsVer07Numerical].
@@
In this section we write $A(u)$ and similar terms in the LHS of an IVP. Later in the text we will resume writing them in the RHS.
Assume $T$ is fixed. Let $h=\frac{T}{N}$ be a fixed time step for some integer $N.$ Set $t_{n}=nh$ so $t_{0},\ldots,t_{N}$ form a partition of $[0;T].$
The following scheme is a general version of the Trotter-Kato formula and is called the Lie-Trotter splitting scheme. In particular, it is a the general version of Theorem 1:
@@
$\frac{dv_{j,n}(t)}{dt}+A_{j}v_{j,n}=0,\quad t\in[t_{n};t_{n+1}],\quad j=\overline{1,m},$
@@
$u(t_{n+1})\approx v_{m,n}(t_{n+1}),\quad n=\overline{0,N-1}.$
@@
$v_{1,n}(t_{n})=\widetilde{v}_{1,n-1}(t_{n}),\ v_{2,n}(t_{n})=v_{1,n}\Big{(}t_{n}+\frac{h}{2}\Big{)},\ \widetilde{v}_{1,n}(t_{n})=v_{2,n}(t_{n+1}),$
$\widetilde{v}_{1,-1}(0)=u_{0}.$
@@
$u(t_{n+1})\approx\widetilde{v}_{1,n}(t_{n+1}),\quad n=\overline{0,N-1}.$
One does expect the second order in the general case but, as earlier, this is not guaranteed.
We will not repeat this universal remark on the absence of a universal rate for any given splitting scheme in future.
@@
The Strang splitting is quite popular not only due to it being of the second order, but also because it does not require doubling the number of computations (compared to the Lie splitting), as we have already seen in the finite-dimensional case. Indeed, if the propagator for the semigroup of $A_{k}$ is $(S^{(k)}_{t})_{t},k=\overline{1,2},$ we get [^21] Here $\prod$ is used instead of $\circ$ to make the expression readable.
$u(T)$ $\approx\widetilde{v}_{2,N-1}(t_{N})$
$=S^{(2)}_{t_{N-1}+h/2,t_{N}}\circ S^{(1)}_{t_{N-1},t_{N}}\circ\prod_{k=\overline{0,N-2}}\Big{[}S^{(2)}_{t_{k}+h/2,t_{k+1}+h/2}\circ S^{(1)}_{t_{k},t_{k+1}}\Big{]}\circ S^{(2)}_{0,h/2}u_{0}.$
The Lie-Trotter and Strang splitting schemes are multiplicative splitting schemes: the formal semigroups for $A_{1},\ldots,A_{m}$ are combined via composition exclusively.
@@
or by using the implicit backward Euler method
@@
The Douglas-Rachford splitting can be extended to the case $m>2$[GlowOshYin17Splitting].
@@
If the operator $A_{2}$ is computationally problematic (e.g. multivalued), both Peaceman-Rachford and Douglas-Rachford splittings can be modified to exclude $A_{2}$ from the second subproblem. For instance, solving the first equation for $A_{2}(\widetilde{t}_{n},\widetilde{u}_{n})$ we get for the Peaceman-Rachford splitting
@@
There are other common additive operator splitting schemes such as the Tseng splitting (e.g. [SuChoGiMouk21parallel, AlGe19iteration]) and the Davis-Yin splitting (e.g. [AraTo22direct, ChenChangLiu20three_operator]).
@@
Let us recall some very basic facts about this last assumption and optimization theory [BauCom17convex, Zeid90nonlinear_2b] and briefly explain some terminology that is often encountered in the context of operator splitting methods in optimization. Let $B$ be a Banach space. A possibly nonlinear and multivalued operator $A\colon B\mapsto B^{*}$ is called monotone if
@@
Operator monotonicity is one of cornerstones of optimization theory in general and convex optimization in particular. The Browder-Minty theorem states that if $A$ is additionally hemicontinuous and coercive [^22] Recall that these are also standard assumptions in the theory of SPDEs., that is, the mappings
@@
Moreover, popular and effective alternating direction methods of (Lagrange) multipliers (ADDM) are actually the reformulated Peaceman-Rachford and Douglas-Rachford splittings (or other additive operator splitting schemes).
@@
([BauCom17convex]) A so-called forward-backward splitting can be derived as follows: we want to find $x$ such that
@@
(5.4) $\frac{du_{h}(t)}{dt}=B_{h}u_{h},$
where the matrix $B$ depends on the grid and the size $h$ of the grid explicitly [^23] If a boundary-value problem is considered, a free term appears in the RHS of the equation.. If both space and time discretizations are used, we have a high-dimensional (possibly nonlinear) equation.